2 log_3(Y)-log_3(y+4) = 2
find y how do you do it ???

Answer by stanbon(75887)

2 log_3(Y)-log_3(y+4) = 2
---
log3(y^2) - log3(y+4) = 2
---
log3[y^2(y+4)] = 2
y^3 + 4y^2 = 3^2
y^3 + 4y^2 - 9 = 0
---
I graphed it and found a root at x=-3
Use synthetic division to find the remaining 2 roots:
---
-3)....1....4....0....-9
........1.....1....-3..|..0
Solve x^2 + x -3 = 0 using the quadratic formula:
x = [-1 +- sqrt(1 -4*1*-3)]/2
x = [-1 +- sqrt(13)]/2
=================================
Check each of the 3 answers to see if any are extraneous.
=================================
Cheers,
Stan H.